OpenStudy (anonymous):

(16+8)×(8+x)/(16+8)+(8+x)=(16×8)/(16+8)+(8×x)/(8+x) What is X? answer is 4 Why answer is 4?

OpenStudy (jdoe0001):

\(\bf \cfrac{(16+8)\times (8+x)}{(16+8)+(8+x)}=\cfrac{16\times 8}{\frac{(16+8)+(8\times x)}{8+x}} ?\)